OpenStudy (anonymous):

x + (x+3) + (2x+6) = 45

OpenStudy (anonymous):

combine like terms x +x +3 +2x +6 = 45 4x +9 = 45 -9 -9 4x = 36 divide both sides by 4 to isolate the x